Tuesday, February 27, 2007

Back to it

Meetings to do with seminar feedback today. Very positive.

Finished off Raman marking, then got some papers off Howarth, including one of the last lot, which has been dealt with, also saw anasuya about her marking, which needs swiftly shifting.

Problem class.

One run finished. Saturn500 finished the 40eV SD-1 run, was analysed recoded as 12eV SD-2 and reset off. This was the penultimate 40eV run remaining and means only two non 12eV runs remain to be completed. Also, only one of the SD 12eV slots is yet to be set off.

Labels: , , , , , , ,

0 Comments:

Post a Comment

<< Home